The Sun River beverage company is a regional producer of teas, exotic juices, and energy drinks. With an interest in healthier lifestyles, there has been an increase in demand for their sugar-free formulation.

The final packing operation requires 13 tasks. Sun River bottles their sugar-free product 5 hours a day, 5 days a week. Each week, there is a demand for 3,000 bottles of this product.

TASK PERFORMANCE TIME (MINUTES) TASK MUST FOLLOW
1 0.1
2 0.1 1
3 0.1 2
4 0.2 2
5 0.1 2
6 0.2 3, 4, 5
7 0.1 1
8 0.1 7
9 0.2 8
10 0.1 9
11 0.2 6
12 0.2 10, 11
13 0.1 12


Using the above data, solve the assembly-line balancing problem and calculate the efficiency of your solution. Use the longest task time for your decision criteria. (Round your answer to the nearest whole percent.)


Efficiency of line balance at cycle time_______ %
Answer #1

Cycle time = 1.8 min.

Takt time = number of minutes available / no of units to be produced

= 5x5x60 /3000 = 0.5 min.

No. of workstations needed = 1.8 /0.5 = 3.6 =4 workstations

2 st tiud Spen 1.8 erriveny